OpenStudy (anonymous):

use the limit definition of definite integral to evalute int_{0}^{1} (2x+3)dx help me please

OpenStudy (saifoo.khan):

\[\int\limits_{0}^{1} (2x+3)dx\]Like this, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and 0 in down

OpenStudy (saifoo.khan):

Now we have to integrate that then install the limits 1 and 0. Agree?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Okay, so integral will be: \[[\frac{2x^2}{2}+ 3x]^1 _0\]

OpenStudy (anonymous):

no please not in this way for rumin sum

OpenStudy (saifoo.khan):

Oh riemann sum.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you for picture alredy i solve it like you and get 4 but in this case want to solve in rieman sum

OpenStudy (anonymous):

please help me i wail waiting

OpenStudy (anonymous):

please ????

OpenStudy (sirm3d):

\[f(x)=2x+3,a=0,b=1,\Delta x = \frac{1}{n}\]

OpenStudy (sirm3d):

\[x_i=a+i \Delta x\]compute \[f(x_i)\]and plug everything into \[\lim_{n \rightarrow +\infty}\sum_{i=1}^{n}f(x_i) \Delta x\]

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

i don't know what you mean for xi

OpenStudy (sirm3d):